One mole of gas having Cp=7/2R id heated at constant pressure from 0°C.What must be the quantity of heat that should be supplied to the gas for the temperature to be doubled?If the same heat is supplied to the gas at constant volume what will be the final temperature?

Grade:11

1 Answers

Harshad aggarwal
6 years ago
Initial temp-273 k final temp-2*271
At constant pressure dq=ncpdt.  dq=1*7/2*R*273
Now at constant volume 
Dq=ncvdt
7/2R*273=5/2R(Tf-273)
On solving Tf=12*273/5